How much energy is required to heat 87.1 g acetone (molar mass=58.08 g/mol) from a solid at -154.0°C to a liquid at -42.0°C? The following physical data may be useful.
1
Expert's answer
2019-08-29T03:39:51-0400

Heating from -154.0 C to melting point (-95 C) (solid state):

"Q1 = C*n*(t2-t1)=96 J\/(mol*K)*10^-3kg\/(0.058kg\/mol)*(-95+154)K=8.5*10^3J"


Melting:

"Q2=\u0394fusH\u2217h=5.72\u221710 \n^3\n J\/mol\u22171.5mol=8.6\u221710 \n^3\n J"


Heating from -95 C to -42 C (liquid state):

"Q3=C\u2217n\u2217(t3\u2212t2)=126.6J\/(mol\u2217K)\u221710 ^\n\u2212\n 3kg\/(0.058kg\/mol)\u2217(\u221242+95)K=10.1\u221710 ^\n3\n J"


The answer:

"Qsum=Q1+Q2+Q3=8.5\u221710 \n^3\n J+8.6\u221710 ^\n3\n J+10.1\u221710 ^\n3\n J=2.7\u221710 ^\n4\n J"